OpenStudy (anonymous):

A 300-page book is 10 inches long, 8 inches wide, and 3 4 inches tall. What is the volume of 100 pages of the book?

RhondaSommer (rhondasommer):

34 inces tall?

OpenStudy (boldjon):

volume = 10 * 8 * .75 = 60 in^3 volume of 100 page book would be 100 / 300 * 60 = 20 in^3

OpenStudy (anonymous):

I feel page count will only change one of the dimensions, A 500 page novel can have the same cover size as a 200 page novel.

OpenStudy (anonymous):

would it be 60 or 20 in^3?